(a) What is the gravitational potential energy of a two-particle system with masses 3.8 kg and 1.1 kg, if they are separated by 1.2 m? If you triple the separation between the particles, how much work is done (b) by the gravitational force between the particles and (c) by you?
Answer #1

a] gravitational potential energy = -Gm1m2/r = -6.67e-11*3.8*1.1/1.2 = -2.32*10^-10 J

b]Work done by gravitational force = - change in PE = 6.67e-11*3.8*1.1*(-1/1.2 + 1/(3*1.2)) = -1.55*10^-10 J

C] WORK done by me = - work done by Gravitational force = 1.55*10^-10 J